Check My Work (No more tries available eBook Current and Quick Ratios Ace Industries has current assets equal to $6 mllion. The companys current ratio is 2.5, and its quick ratio is 2.0. Do not answers to the nearest dollar What is the firms level of current liabilities? 2400000 What is the firms level of inventories? 1200000
Answer #1
a)
we know that current ratio = Current asset / Current liabilities
Therefore
Current asset / Current liabilities = 2.5
Current liabilities = 6000000/2.5
Current liabilities 2400000
b)
Given that quick ratio = 2
this means
(Current asset - inventory)/ Current liabilities = 2
=(6000000-inventory)/2400000= 2
inventory = 6000000-4800000
inventory = 1200000